Questions 1: Consider the titration of a 24.0-mL sample of 0.175 M CH3NH2 with 0.155 M HBr. (The value of Kb for CH3NH2 is 4.4×10−4

A) Determine the initial pH

B) Determine the volume of added acid required to reach the equivalence point

C) Determine the pH at 4.0 mL of added acid

D) Determine the pH at one-half of the equivalence point.

E) Determine the pH at the equivalence point.

F) Determine the pH after adding 5.0 mL of acid beyond the equivalence point.
Answer #1

CH3NH2 = 24.0mL of 0.175M

HBr= 0.155M

Kb = 4.4x10^-4

a) Initial PH

Concentration of CH3NH2 = C= 0.175M

for weak bases

[OH-] = square root of Kb xC

[OH-]= square root of ( 4.4x10^-4 x 0.175)

[OH-]= 0.877 x10^-2

-log[OH-]= -log[0.877x10^-2]

POH = 2.06

PH + POH = 14

PH = 14 - POH

PH = 14 - 2.06

PH = 11.94

b) volume at equivalent

volume of HBr at equivalent point = 0.175 x 24.0 / 0.155 = 27.1 mL

volume of HBr = 27.1 mL

c) at the addition of 4 mL

CH3NH2 = 24.0mL of 0.175M

number of moles of CH3NH2 = 0.175M x 0.0240L = 0.0042 moles

HBr = 4.0mL of 0.155M

number of moles of HBr = 0.155M x 0.004L = 0.00062 moles

Kb = 4.4x10^-4

-log(Kb) = -log( 4.4x10^-4)

PKb = 3.36

           CH3NH2 +   HBr ----------------- CH3NH3+Br-

          0.0042           0.00062                  0

      - 0.00062         - 0.00062              + 0.00062

        0.00358                  0                  + 0.00062

POH = Pkb + log[salt]/[base]

POH = 3.36 + log( 0.00062/0.00358)

POH = 2.60

PH = 14 - POH

PH = 14 - 2.60

PH = 11.4

d) at one-half equivalent point

POH = PKb

POH = 3.36

PH = 14 - 3.36 = 10.64

PH = 10.64

e) at equivalent point

number of moles of CH3NH2 = 0.175M x 0.0240L = 0.0042 moles

HBr= 27.1mL of 0.155 moles

number of moles of HBr = 0.155M x 0.0271L = 0.0042 moles

at equivelent point, number of moles of CH3NH2 is equal to HBr

Total volume = 24.0+27.1 = 51.1ml = 0.0511L

Concentration = C = number of moles /Total volume = 0.0042/0.0511 = 0.0822 M

at equivalent point

PH = 7 - 1/2[PKb + logC]

PH = 7 - 1/2[3.36+ logC]

PH = 7 - 1/2[3.36 + log(0.0822)]

PH = 5.86

f) after adding 5 mL of acid beyond equivalent point

volume of Acid = 27.1 + 5.0 = 32.1 ml

number of moles of acid = 0.155M x 0.0321L = 0.00497 moles

number of moles of CH3NH2 = 0.175M x 0.0240L = 0.0042 moles

number of moles of acid is greater than the base

so the solution is acidic nature

remaining number of moles of acid = 0.00497 - 0.0042 = 0.00077 moles

Total volume = 24.0 + 32.1 = 56.1mL = 0.0561L

Concentration of H+ = number of moles/volume = 0.00077 / 0.0561 = 0.0137 M

[H+] = 0.0137M

-log[H+] = -log( 0.0137)

PH = 1.86.